OpenStudy (anonymous):

x^2+y^2+4x-6y=12 find the radius and the center of the cirlce

OpenStudy (anonymous):

center(-2,3) radius=sqrt(4+9+12)=5

OpenStudy (anonymous):

Groups it like this: (x^2+4x) + (y^2+6y) = 12 Complete the square to get: (x+2)^2 + (y+3)^2 = 25 sqrt(25) = 5

OpenStudy (anonymous):

y^2-6y